# Genesis-Metallicity: Universal Non-Parametric Gas-Phase Metallicity Estimation

### v2

@@ 5 Conclusion
Improved sampling of the sparsely populated regions of the emission line observables parameter space can further enhance the accuracy of our calibration. Therefore, we commit to keeping genesis-metallicity and its calibration data up-to-date in light of the upcoming data. The most recent version of genesis-metallicity and its calibration data can be found at https://github.com/langeroodi/genesis_metallicity (the calibration data will be made available upon publication). The version corresponding to this draft (1.2.0) is archived at Zenodo and can be found at Langeroodi (2025).
@@ Acknowledgments
We greatly appreciate comments from the referee on both the manuscript and the public code, which substantially improved the presentation of the data and methodology, as well as the accessibility of the software. We are grateful to the statistics editor, whose feedback helped shape the discussion in Appendix C. We also appreciate comments from Chris Willott, which were crucial in identifying a bug in the initial version (v1.0) of the public code. This work was made possible by the public release of the reduced JWST NIRSpec MSA spectra acquired through the JADES and JOF programs. Moreover, we heavily used the latest release of the SDSS data (DR18). This work was supported by research grants (VIL16599, VIL54489) from VILLUM FONDEN.
@@ Appendix B Literature Parametric Calibrations
Parametric strong-line calibrations often rely on polynomial fits to projections of calibration data in line ratio vs. metallicity 2D planes. As discussed in Section 1, these calibrations risk overlooking the complexities of the higher-order parameter space. Most often, even the 2D projections cannot be fully captured by polynomials. This is evident from the large scatter of calibration data around the best-fit polynomials in 2D projection planes. This has resulted in large discrepancies between different strong-line calibrations reported in the literature. We show this in Figure 15, where some widely adopted parametric strong-line calibrations are compared with our calibration data.
@@ Appendix C Multivariate KDE implementation
In this Section, we discuss the details of our multivariate KDE implementations presented in Sections 3.1 and 4. Our KDE bandwidth selection is described in Section C.1, while reliability tests beyond those presented in the main text are provided in Section C.2.
@@ C.1 Bandwidths
In this work, we adopt Gaussian kernels with bandwidths selected through a cross-validation algorithm. Bandwidth selection is performed using the statsmodels library (Seabold & Perktold, 2010), adopting the maximum likelihood cross-validation approach suggested in Li & Racine (2007) which was originally proposed in Duin (1976). The selected bandwidths are then imported into our scipy implementation of KDE, described in Sections 3.1 and 4. The selected bandwidths for our (O II) electron temperature estimator (Section 3.1) in the 5-dimensional space of O2, O3, EW(H), (O II), and (O III) are reported in Table 2. Moreover, the selected bandwidths for our gas-phase metallicity estimator (Section 4) in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity are reported in Table 3.
@@ C.2 Validation
As the dimensionality of the parameter space increases, multivariate KDEs should be used with caution because the curse of dimensionality leads to an exponential growth in the sample size required to maintain constant accuracy (Scott, 1992; Di Marzio & Lafratta, 1999; Nagler & Czado, 2015). This stems from an exponential increase in the variance term as the number of KDE covariates grows. Therefore, the addition of each new covariate must be carefully weighed against the trade-off between information gain and the variance inflation induced by the curse of dimensionality. As discussed in Sections 3.1 and 4, by removing the EW(H) covariate we construct lower-dimensional KDEs for both our electron temperature and gas-phase metallicity estimators, finding a decline in the accuracy of both. This indicates that by including the EW(H) axis we are winning the trade-off between information gain and variance inflation in both cases.

### v3

@@ V Conclusion
Improved sampling of the sparsely populated regions of the emission line observables parameter space can further enhance the accuracy of our calibration. Therefore, we commit to keeping genesis-metallicity and its calibration data up-to-date in light of the upcoming data. The most recent version of genesis-metallicity and its calibration data can be found at https://github.com/langeroodi/genesis_metallicity. The version corresponding to this draft (1.2.0) is archived at Zenodo and can be found at Langeroodi (2025).
@@ Acknowledgments
We greatly appreciate comments from the referee on both the manuscript and the public code, which substantially improved the presentation of the data and methodology, as well as the accessibility of the software. We are grateful to the statistics editor, whose feedback helped shape the discussion in Appendix C. We also appreciate comments from Chris Willott, which were crucial in identifying a bug in the initial version (v1.0) of the public code. This work was made possible by the public release of the reduced JWST NIRSpec MSA spectra acquired through the JADES and JOF programs; all the JWST data used in this work can be found in MAST: http://dx.doi.org/10.17909/8tdj-8n28 (catalog 10.17909/8tdj-8n28). Moreover, we heavily used the latest release of the SDSS data (DR18). This work was supported by research grants (VIL16599, VIL54489) from VILLUM FONDEN.
@@ Appendix B Literature Parametric Calibrations
Parametric strong-line calibrations often rely on polynomial fits to projections of calibration data in line ratio vs. metallicity 2D planes. As discussed in Section I, these calibrations risk overlooking the complexities of the higher-order parameter space. Most often, even the 2D projections cannot be fully captured by polynomials. This is evident from the large scatter of calibration data around the best-fit polynomials in 2D projection planes. This has resulted in large discrepancies between different strong-line calibrations reported in the literature. We show this in Figure 15, where some widely adopted parametric strong-line calibrations are compared with our calibration data.
@@ Appendix C Multivariate KDE implementation
In this Section, we discuss the details of our multivariate KDE implementations presented in Sections III.1 and IV. Our KDE bandwidth selection is described in Section C.1, while reliability tests beyond those presented in the main text are provided in Section C.2.
@@ C.1 Bandwidths
In this work, we adopt Gaussian kernels with bandwidths selected through a cross-validation algorithm. Bandwidth selection is performed using the statsmodels library (Seabold and Perktold, 2010), adopting the maximum likelihood cross-validation approach suggested in Li and Racine (2007) which was originally proposed in Duin (1976). The selected bandwidths are then imported into our scipy implementation of KDE, described in Sections III.1 and IV. The selected bandwidths for our (O II) electron temperature estimator (Section III.1) in the 5-dimensional space of O2, O3, EW(H), (O II), and (O III) are reported in Table 2. Moreover, the selected bandwidths for our gas-phase metallicity estimator (Section IV) in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity are reported in Table 3.
@@ C.2 Validation
As the dimensionality of the parameter space increases, multivariate KDEs should be used with caution because the curse of dimensionality leads to an exponential growth in the sample size required to maintain constant accuracy (Scott, 1992; Di Marzio and Lafratta, 1999; Nagler and Czado, 2015). This stems from an exponential increase in the variance term as the number of KDE covariates grows. Therefore, the addition of each new covariate must be carefully weighed against the trade-off between information gain and the variance inflation induced by the curse of dimensionality. As discussed in Sections III.1 and IV, by removing the EW(H) covariate we construct lower-dimensional KDEs for both our electron temperature and gas-phase metallicity estimators, finding a decline in the accuracy of both. This indicates that by including the EW(H) axis we are winning the trade-off between information gain and variance inflation in both cases.